# RING-NeRF: A Versatile Architecture based on Residual Implicit Neural Grids

### v1

@@ 9.1 Choice of the baselines and Implementation
While NeuS-facto is solely implemented in the sdfstudio [20] framework (which derives from nerfstudio), Neuralangelo’s code has been officially released and also possesses a SdfStudio implementation. It is to be noted that, while the specificities of the article are coded in a similar fashion, both implementations make many different choices on more general NeRF subjects. For instance, a major difference between both frameworks is the way they sample rays at each iteration. While SdfStudio randomly chooses pixels among all the images (as in Nerfstudio), Neuralangelo follows I-NGP framework by sampling every pixels of $n$ images ($n$ being the batch size). Both solutions are valid but they enable different behaviors. In order to harmonize as much as possible our different experiments, we decided to use the SdfStudio implementation, as the major part of the framework is common with NerfStudio, which was used in the other experiments.
@@ 9.2 Configuration Details
Our experiments were all done on the SdfStudio framework, using the provided replica subset dataset. We use all the scenes, except the scan5, for which we experienced failure of training for every method without initialization, and expect the darker environment to heavily complexify the reconstruction in an already hard initialization-lacking setup. The estimated depths are provided by SdfStudio and predicted via a pre-trained OmniData [^6] A. Eftekhar, A. Sax, J. Malik, and A. Zamir, Omnidata: A scalable pipeline for making multi-task mid-level vision datasets from 3d scans. In Proc. of the IEEE International Conf. on Computer Vision (ICCV), 2021 model, they serve as an indication rather than true ground truths.
We mostly follow the basic Neuralangelo configuration implemented by SdfStudio. This results in 16 levels of resolutions 32 to 4096 with features of size 8 and a hashmap size of $2^{19}$. The main differences regards the scheduling of the training, initially implemented to last 500k epochs. However, we noticed a similar convergence when scaling down the scheduling parameters for the training to last 100k epochs. Hence, we decided to evaluate all three methods (RING-NeRF, NeuS-facto and Neuralangelo) on both 20K epochs training and 100K epochs (with accordingly scaled down schedulers and coarse-to-fine duration). Because the replica dataset is a synthetic inside dataset, we remove background MLPs from all evaluated baselines, appearance embedding and scene contraction.
Note that differently from the two other baselines, RING-NeRF does not input the position on its MLP, which results in a different way to code the initialization. NeuS-facto and Neuralangelo both hard code the SDF inverted sphere in the weights of the MLP, benefiting from the position input while we initialize the model randomly and then overfit the network for 1000 epochs to the required initialization scheme before beginning the ”true” training on the scene.
@@ 9.3 Results
Additional results of SDF training without initialization on several Replica scenes can be found in figure 12. While most training of NeuS-facto without initialization results in catastrophic failure, some scenes still succeed in producing a coherent reconstruction. The scan 6, shown in figure 12, is one of those scenes, and while NeuS-facto avoids failure, the results are still way noisier than our RING-NeRF both in RGB and depth, and results in way lower PSNR metrics, both for 20k and 100k configurations. Similarly, RING-NeRF without coarse-to-fine will often face failure, and therefor needs its continuous coarse-to-fine to forego the initialization.
Results with initialization. While not the subject of the initial experiment, it is interesting to compare those results with their correctly initialized counterparts, as well as in-between initialized models. Figure 9 gives an example on one scene of the dataset.
First of all, our approach presents similar results whether we initialize the model or not. This further proves that RING-NeRF is independent of the initialization process. We can notice a slight difference in the PSNR metric which is within the randomness interval.
@@
Regarding Neuralangelo, initializing the model improves both the RGB metrics and the depth estimation. However, the depth is still surprisingly blurry, which in turns prevents the RGB renders to overcome RING-NeRF results. It is however important to note that combining Neuralangelo’s numerical gradients and curvature loss with RING-NeRF architecture creates satisfying renders which are close to the sole RING-NeRF. This further seems to indicate that, in addition to improving the robustness to initialization, our RING-NeRF architecture enables better depth estimation by design.
@@
10.1 Configuration Details
Our pre-training was done on $N=10$ scenes chosen at random of the DTU dataset (the IDs used for the article’s experiment are : 24, 63, 65, 69, 83, 97, 105, 106, 118, 122). It is done with a similar pipeline as in Control-NeRF [9]. We create $N$ sets of hashgrids, one for each scene and one unique decoder (both density and color networks). For each epoch, we randomly create a batch of training rays chosen among all the images of all the scenes and then dispatch them properly among the corresponding grids.
Since the decoder latent space is invariant to position and LOD, a pre-trained decoder can be used for different grid resolutions and hierarchy size. The hierarchy size can even be adjusted dynamically during the reconstruction.
In order to accentuate the difference when adding a new grid, we used a very simple base configuration for this experiment with a HashGrid pyramid of 3 levels of respective resolutions of 16, 32 and 64 and once again used features of size 8. While the dictionary of size $2^{19}$ is big enough to have the equivalent of dense features for this configuration, we still use HashGrids as we then proceed to increase the resolution up to a higher number of parameters than the hashmap size.
10.2 Results and Utility of the property
Figure 10 shows novel view synthesis of one scene both when training from scratch and with the decoder pre-trained and frozen. As shown in the article (table 5), the models trained from scratch gives slightly better results. However, the use of a frozen pre-trained decoder enables extensibility in resolution. While this property can seem unimportant at first, it actually is crucial for embedded systems with important resources limitations. We can for instance imagine two use cases of this property :
  • A SLAM situation, where the robot is moving without interruption. Because it is constantly discovering new environment to reconstruct, it does not have enough time to reconstruct with precision the previously seen areas. However, RING-NeRF architecture will let the robot optimize with higher resolution previous areas later on when it will have more free resources.
  • An important issue in grid-based NeRFs is the choice of hyper-parameters, and more especially the maximum resolution of the grid pyramid. Being able to dynamically change the resolution of your model gives the possibility to dynamically choose the optimal maximum resolution during training depending on the scene’s complexity, therefore discarding any useless parameters for a more optimal model.
Figure 10: Resolution extensibility example. PSNR of the corresponding image are inset. ”Free” means that the decoder is learnt jointly with the grids from scratch and ”Frozen” means that the decoder is frozen and pre-trained and only the grids are trained. ”+1 LvL” also means that we use the frozen checkpoint of the Frozen 3Lvls experiment and only train the last set of grid.
@@ 10.2 Results and Utility of the property
PSNR ↑ Outside Inside
Scene Bicycle Flowers Garden Stump Treehill Room Counter Kitchen Bonsai
NeRF 21.76 19.40 23.11 21.73 21.28 28.56 25.67 26.31 26.81
Mip-NeRF 21.69 19.31 23.16 23.10 21.21 28.73 25.59 26.47 27.13
Mip-NeRF 360 24.40 21.64 26.94 26.36 22.81 31.40 29.44 32.02 33.11
Nerfacto 24.75 21.43 27.05 26.11 22.71 31.56 26.33 30.54 33.37
Zip-NeRF (repr.) 25.19 21.84 27.63 26.84 23.42 32.42 28.49 31.76 34.96
Our Model 25.29 21.91 27.98 26.70 23.18 32.24 28.95 31.92 34.80
Table 8: Novel View Synthesis PSNR for the Mono-Scale setup per scene on the 360 Dataset.
SSIM ↑ Outside Inside
Scene Bicycle Flowers Garden Stump Treehill Room Counter Kitchen Bonsai
NeRF 0.455 0.376 0.546 0.453 0.459 0.843 0.775 0.749 0.792
Mip-NeRF 0.454 0.373 0.543 0.517 0.466 0.851 0.779 0.745 0.818
Mip-NeRF 360 0.693 0.583 0.816 0.746 0.632 0.913 0.895 0.920 0.939
Nerfacto 0.694 0.579 0.819 0.730 0.590 0.911 0.844 0.907 0.940
Zip-NeRF (repr.) 0.730 0.604 0.845 0.763 0.636 0.929 0.885 0.922 0.961
Our Model 0.724 0.606 0.848 0.760 0.613 0.920 0.876 0.917 0.951
Table 9: Novel View Synthesis SSIM for the Mono-Scale setup per scene on the 360 Dataset.
LPIPS ↓ Outside Inside
Scene Bicycle Flowers Garden Stump Treehill Room Counter Kitchen Bonsai
NeRF 0.536 0.529 0.415 0.551 0.546 0.353 0.394 0.335 0.398
Mip-NeRF 0.541 0.535 0.422 0.490 0.538 0.346 0.390 0.336 0.370
Mip-NeRF 360 0.289 0.345 0.164 0.254 0.338 0.211 0.203 0.126 0.177
Nerfacto 0.212 0.259 0.121 0.192 0.372 0.121 0.187 0.0934 0.0709
Zip-NeRF (repr.) 0.186 0.246 0.105 0.151 0.297 0.111 0.146 0.0812 0.0592
Our Model 0.182 0.232 0.0908 0.167 0.309 0.118 0.148 0.0851 0.0598
Table 10: Novel View Synthesis LPIPS for the Mono-Scale setup per scene on the 360 Dataset.
@@
Mip-NeRF 360 24.51 26.93 28.53 29.24 21.64 23.90 26.01 27.35 26.71 29.59 31.35 32.52 26.27 27.68 28.82 29.27 22.93 24.63 26.06 27.12
Zip-NeRF (repr.) 24.85 27.71 29.08 29.99 21.51 25.06 27.37 28.61 27.03 29.83 31.38 31.95 25.94 27.83 29.24 30.25 22.93 25.55 27.36 28.56
Our Model 25.02 27.87 29.31 29.92 21.80 25.38 27.34 28.53 27.52 30.27 31.60 32.09 26.08 27.81 29.28 30.29 23.25 26.00 27.83 29.01
@@
Mip-NeRF 360 31.44 32.53 33.17 32.96 29.30 30.12 30.81 30.52 31.90 33.39 34.69 34.92 32.85 33.97 34.63 33.80
Zip-NeRF (repr.) 31.65 32.52 32.78 .74 28.19 29.15 29.45 29.49 31.72 32.21 32.51 32.53 34.07 34.20 34.03 33.94
Our Model 31.53 32.45 32.55 32.30 28.35 29.22 29.72 29.78 31.67 32.25 32.68 32.21 33.87 34.29 33.97 33.55
@@
Mip-NeRF 360 0.666 0.815 0.890 0.912 0.567 0.727 0.834 0.881 0.791 0.903 0.939 0.959 0.726 0.819 0.874 0.882 0.615 0.748 0.839 0.893
Zip-NeRF (repr.) 0.703 0.834 0.887 0.896 0.565 0.752 0.835 0.866 0.820 0.897 0.923 0.926 0.725 0.807 0.847 0.845 0.591 0.758 0.831 0.874
Our Model 0.704 0.839 0.888 0.894 0.574 0.761 0.835 0.861 0.820 0.891 0.920 0.927 0.722 0.804 0.843 0.848 0.600 0.766 0.850 0.892
@@
Mip-NeRF 360 0.906 0.944 0.963 0.967 0.887 0.916 0.936 0.942 0.916 0.949 0.968 0.975 0.935 0.959 0.969 0.968
Zip-NeRF (repr.) 0.910 0.932 0.937 0.936 0.857 0.869 0.875 0.881 0.909 0.924 0.932 0.941 0.946 0.945 0.938 0.939
Our Model 0.901 0.925 0.931 0.930 0.851 0.875 0.891 0.900 0.905 0.917 0.924 0.932 0.939 0.944 0.938 0.941
@@
Mip-NeRF 360 0.322 0.177 0.089 0.066 0.367 0.215 0.114 0.071 0.194 0.079 0.045 0.029 0.279 0.171 0.114 0.107 0.362 0.236 0.144 0.096
Zip-NeRF (repr.) 0.180 0.106 0.0622 0.0428 0.306 0.113 0.0639 0.0471 0.135 0.0637 0.0436 0.0358 0.203 0.121 0.0959 0.0719 0.390 0.191 0.119 0.0709
Our Model 0.218 0.0932 0.0535 0.0468 0.284 0.104 0.0701 0.0725 0.119 0.0539 0.0423 0.0369 0.227 0.129 0.104 0.0850 0.355 0.171 0.0978 0.0601
@@
Mip-NeRF 360 0.227 0.101 0.052 0.042 0.216 0.114 0.068 0.059 0.134 0.063 0.033 0.023 0.185 0.065 0.033 0.033
Nerfacto 0.320 0.163 0.0714 0.0931 0.438 0.288 0.166 0.176 0.258 0.113 0.0918 0.176 0.233 0.112 0.0853 0.160
Zip-NeRF (repr.) 0.141 0.0644 0.0355 0.0241 0.188 0.118 0.0642 0.0612 0.0977 0.0636 0.0426 0.0286 0.0787 0.0503 0.0454 0.0373
Our Model 0.161 0.0692 0.0362 0.0263 0.204 0.0992 0.0553 0.0395 0.107 0.0596 0.0410 0.0356 0.0864 0.0477 0.0445 0.0369

### v3

@@ 9.1 Choice of the baselines and Implementation
While NeuS-facto is solely implemented in the Sdfstudio [26] framework (which derives from Nerfstudio), Neuralangelo’s code has been officially released and also possesses a Sdfstudio implementation. It is to be noted that, while the specificities of the article are coded in a similar fashion, both implementations make many different choices on NeRF specificities. For instance, a major difference between both frameworks is the way they sample rays at each iteration. While SdfStudio randomly chooses pixels among all the images (as in Nerfstudio), Neuralangelo follows I-NGP framework by sampling every pixels of $n$ images ($n$ being the batch size). Both solutions are valid but they enable different behaviors. In order to harmonize as much as possible our different experiments, we decided to use the Sdfstudio implementation, as the major part of the framework is common with Nerfstudio, which was used in the other experiments.
@@ 9.2 Configuration Details
Our experiments are all done on the Sdfstudio framework, using the provided replica subset dataset. We use all the scenes, except the scan5, for which we experienced failure of training for every method without initialization, and expect the darker environment to heavily complexify the reconstruction in an already hard initialization-lacking setup. The estimated depths are provided by Sdfstudio and predicted via a pre-trained OmniData [^5] A. Eftekhar, A. Sax, J. Malik, and A. Zamir, Omnidata: A scalable pipeline for making multi-task mid-level vision datasets from 3d scans. In Proc. of the IEEE International Conf. on Computer Vision (ICCV), 2021 model, they serve as an indication rather than true ground truths.
We mostly follow the basic Neuralangelo configuration implemented by Sdfstudio. This results in 16 levels of resolutions 32 to 4096 with features of size 8 and a hashmap size of $2^{19}$. The main differences regards the scheduling of the training, initially implemented to last 500k epochs. However, we noticed a similar convergence when scaling down the scheduling parameters for the training to last 100k epochs. Hence, we decided to evaluate all three methods (RING-NeRF, NeuS-facto and Neuralangelo) on 100K epochs (with accordingly scaled down schedulers and coarse-to-fine duration). Because the replica dataset is a synthetic indoor dataset, we remove background MLPs, appearance embedding and scene contraction from all evaluated baselines.
In the following additional experiments, we provide more results without initialization, as well as results with correct scene-specific initialization of the scene using SDF inverted spheres. Note that because RING-NeRF does not input the position on its MLP, it results in a different way to code the initialization. NeuS-facto and Neuralangelo both hard code the SDF inverted sphere in the weights of the MLP, benefiting from the position input while we initialize the model randomly and then overfit the network for 1000 epochs to the required initialization scheme (inverted sphere) before beginning the "true" training on the scene.
@@ 9.3 Results
Additional results of SDF training without initialization on several Replica scenes can be found in figure 19. While most training of NeuS-facto without initialization results in catastrophic failure, some scenes still succeed in producing a coherent reconstruction. The scan 6, shown in figure 19, is one of those scenes, and while NeuS-facto avoids failure, the results are still way noisier than our RING-NeRF both in RGB and depth, and results in way lower PSNR metrics. Similarly, RING-NeRF without coarse-to-fine will often face failure, and therefore needs its continuous coarse-to-fine mechanism to forego the initialization.
Reconstruction Results. Figure 14 shows some meshes obtained from the models with random initialization. As expected, the NeuS-facto mesh is completely useless, with the cloudy artefacts from the model resulting in opaque matter. On the other hand, both NeuralAngelo and RING-NeRF produces coherent meshes. However, the NeuralAngelo mesh is very blurry and contains very little details, while ours is way more precise, although rather noisy. These qualitative results are coherent with the previously described results.
Figure 14: Resulting Meshes for models without initialization. Chamfer Distances in cm are inset.
Results with initialization. While not the subject of the initial experiment, it is interesting to compare those results with their correctly initialized counterparts. Figure 15 gives an example on one scene of the dataset.
First of all, our approach presents similar results whether we initialize the model or not. This further proves that RING-NeRF is robust to the initialization process. We can notice a slight difference in the PSNR metric which can be considered within the randomness interval of two reconstruction reruns of the same model and scene.
@@
Regarding Neuralangelo, initializing the model improves both the RGB metrics and the depth estimation. However, the depth is still surprisingly blurry, which in turns prevents the RGB renders to overcome RING-NeRF results. This may be caused by the curvature loss, which is supposed to smooth the reconstruction. Because RING-NeRF introduces a bit of noise in the reconstruction, combining thoses two contributions could be an interesting research focus for robust and smooth SDF reconstruction.
@@
Results on real data. We demonstrated that the scene-specific initialization was crucial for other methods on the Replica dataset, contrary to our RING-NeRF. Replica being a perfect synthetic dataset, this further highlights the importance of the initialization in SDF representation. However, we also performed this no-initialization experiment on more complex real data such as the Tanks and Temples dataset. Figure 16 shows results on one such scene ("Meeting Room"). As expected, while NeuS-Facto faces catastrophic failure, both NeuralAngelo and RING-NeRF succeeds in reconstructing the scene, although with varying issues. NeuralAngelo is both blurrier and with more coarse errors (eg. the ceiling’s beams) while RING-NeRF faces some noise. The latter however presents better global results as confirmed by the PSNR values which are even out on a test images subset of the scene.
Figure 16: Depth prediction of SDF reconstructions without initialization on one Tanks and Temple scene. Note that NeuS-Facto fails to reconstruct the scene. PSNR mean values on the test set are inset.
@@
10.1 Demontration of the resolution extensibility on an unbounded complex scene.
Since the resolution extensibility is an intrinsic property of RING-NeRF, the experiment can be reproduced on any scene or configuration. Here, we demonstrate the property on a more complex scene of the 360 dataset, namely Garden. We follow the same experimental protocol described in section 4.5 with an increased hierarchy (6 levels from 16 to 512 max resolution). We begin the reconstruction with only the first three levels, train both the grid and the decoder to convergence and then freeze both of them. We then proceed to train to convergence one novel grid at a time, freezing the previous one at convergence. Figure 17 shows the final reconstruction at different output of the grid hierarchy. We notice that adding a grid always improves the reconstruction (with an increasing PSNR). This showcases both the capacity of our model to reconstruct finer details after the decoder’s training and its ability to keep the coarser LOD valid.
Figure 17: Resolution Extensibility on the Garden scene of the 360 dataset. Apart from the first grid, each novel grid is trained alone, with both the previous grids and the decoder frozen.
10.2 Discussion on the Utility of the resolution extensibility.
While resolution extensibility has not been heavily studied in previous works, it actually is a crucial property in several different use cases and applications. On one side, this can be useful for better compression of NeRF models. An important issue in grid-based NeRFs is the choice of hyper-parameters, and more especially the maximum resolution of the grid pyramid. Being able to dynamically change the resolution of your model gives the possibility to dynamically choose the optimal local maximum resolution during training depending on the scene’s complexity, therefore discarding any useless parameters for a more optimal model. On the other side, the usefulness becomes obligation for embedded systems with important resources limitations. We can for instance imagine two types of limitations in SLAM situations, where the robot is moving without interruption:
  • Time limitation : Because it is constantly discovering new environment to reconstruct, it does not have enough time to reconstruct with precision the previously seen areas. However, RING-NeRF will let the robot optimizes with higher resolution previous areas later on when it will have more resources available.
  • Memory Limitation : As the scene to reconstruct is of unknown size and the system possesses limited memory, it can not afford to reconstruct the entirety of the scene with maximum precision. A more viable strategy would be to reconstruct coarsely the majority of the scene and only reconstruct finer details on important areas (whether the most looked at places or those with the most objects for instance).
@@
Table 8: Novel View Synthesis PSNR for the Mono-Scale setup per scene on the 360 Dataset (outside scenes).
@@
Nerfacto 24.75 25.84 23.84 22.08 21.43 22.54 21.99 20.26 27.05 26.41 23.91 22.20 26.11 23.37 23.38 22.90 22.71 24.99 24.58 23.41
PyNeRF 25.46 22.95 22.37 21.25 21.93 17.73 16.65 16.55 28.55 22.96 21.75 20.71 26.40 21.65 21.40 20.87 23.28 18.50 17.89 17.60
Zip-NeRF 25.19 15.09 14.48 12.55 21.84 14.96 9.59 7.58 27.63 14.86 15.31 13.99 26.84 19.09 13.19 9.67 23.42 17.42 11.88 9.67
Our Model 25.26 26.11 25.63 24.29 21.91 23.35 22.89 21.30 28.03 27.14 25.32 23.50 26.69 26.20 25.82 24.59 23.12 24.93 25.19 23.92
Table 9: Novel View Synthesis PSNR for the Mono-Scale setup per scene on the 360 Dataset (inside scenes).
@@
Nerfacto 31.59 29.77 29.44 26.58 26.33 27.18 25.79 23.81 30.54 29.70 25.95 23.46 33.37 31.45 28.23 25.24
PyNeRF 32.12 24.51 23.52 22.63 27.76 19.38 18.76 18.52 32.56 25.35 24.25 22.86 32.80 24.41 23.73 22.48
Zip-NeRF 32.42 17.07 10.56 8.36 28.49 17.12 10.79 8.19 31.76 19.58 14.45 12.70 34.96 14.01 6.64 4.20
Our Model 32.43 28.02 26.00 24.29 29.06 27.73 25.27 24.36 32.66 29.42 27.42 26.88 35.15 31.68 28.83 26.32
Table 10: Novel View Synthesis SSIM for the Mono-Scale setup per scene on the 360 Dataset (outside scenes).
@@
Nerfacto 0.694 0.774 0.662 0.560 0.579 0.661 0.645 0.551 0.819 0.767 0.632 0.554 0.730 0.613 0.555 0.481 0.590 0.698 0.683 0.647
PyNeRF 0.726 0.734 0.682 0.613 0.592 0.535 0.538 0.543 0.847 0.748 0.638 0.582 0.752 0.637 0.598 0.533 0.632 0.602 0.599 0.586
Zip-NeRF 0.730 0.237 0.156 0.073 0.604 0.581 0.451 0.364 0.845 0.319 0.303 0.251 0.763 0.687 0.471 0.259 0.636 0.654 0.495 0.408
Our Model 0.726 0.768 0.786 0.738 0.601 0.672 0.683 0.628 0.852 0.817 0.751 0.681 0.759 0.731 0.717 0.646 0.610 0.692 0.761 0.755
Table 11: Novel View Synthesis SSIM for the Mono-Scale setup per scene on the 360 Dataset (inside scenes).
@@
Nerfacto 0.911 0.874 0.876 0.827 0.844 0.861 0.808 0.752 0.907 0.869 0.706 0.606 0.940 0.922 0.841 0.749
PyNeRF 0.919 0.827 0.780 0.747 0.873 0.649 0.585 0.567 0.930 0.837 0.745 0.677 0.946 0.852 0.796 0.727
ZipNeRF 0.929 0.769 0.543 0.438 0.885 0.723 0.512 0.406 0.922 0.805 0.637 0.519 0.961 0.593 0.251 0.189
Our Model 0.921 0.899 0.870 0.836 0.876 0.859 0.821 0.765 0.926 0.871 0.808 0.756 0.954 0.928 0.879 0.830
Table 12: Novel View Synthesis LPIPS for the Mono-Scale setup per scene on the 360 Dataset (outside scenes).
@@
Nerfacto 0.212 0.167 0.281 0.331 0.259 0.191 0.259 0.332 0.121 0.183 0.291 0.338 0.192 0.270 0.370 0.471 0.372 0.265 0.283 0.306
PyNeRF 0.196 0.176 0.236 0.287 0.282 0.278 0.341 0.341 0.096 0.183 0.288 0.349 0.160 0.243 0.334 0.480 0.305 0.299 0.310 0.320
Zip-NeRF 0.186 0.651 0.641 0.698 0.246 0.244 0.432 0.508 0.105 0.744 0.645 0.666 0.151 0.207 0.349 0.482 0.297 0.280 0.368 0.422
Our Model 0.184 0.161 0.141 0.164 0.236 0.166 0.159 0.209 0.090 0.112 0.159 0.197 0.168 0.193 0.218 0.246 0.308 0.258 0.187 0.199
Table 13: Novel View Synthesis LPIPS for the Mono-Scale setup per scene on the 360 Dataset (inside scenes).
@@
Nerfacto 0.121 0.091 0.120 0.153 0.187 0.122 0.161 0.207 0.093 0.102 0.198 0.294 0.071 0.069 0.145 0.239
PyNeRF 0.111 0.194 0.208 0.207 0.149 0.293 0.309 0.322 0.076 0.123 0.182 0.262 0.069 0.114 0.166 0.245
Zip-NeRF 0.111 0.215 0.373 0.405 0.146 0.214 0.307 0.351 0.081 0.163 0.353 0.456 0.059 0.380 0.717 0.716
Our Model 0.120 0.087 0.082 0.086 0.150 0.109 0.105 0.119 0.080 0.094 0.120 0.155 0.056 0.062 0.109 0.130
PSNR ↑ Bicycle Flowers Garden Stump Treehill
Res x1 x2 x4 x8 x1 x2 x4 x8 x1 x2 x4 x8 x1 x2 x4 x8 x1 x2 x4 x8
PyNeRF 25.77 28.44 30.12 30.11 22.13 25.80 27.28 27.80 28.35 31.23 32.02 31.54 27.21 28.34 29.87 30.05 23.23 26.17 27.98 29.43
Zip-NeRF 24.85 27.71 29.08 29.99 21.51 25.06 27.37 28.61 27.03 29.83 31.38 31.95 25.94 27.83 29.24 30.25 22.93 25.55 27.36 28.56
Our Model 25.17 27.94 29.21 29.91 21.72 25.27 27.20 28.20 27.40 30.00 31.22 31.70 25.97 27.81 29.05 29.81 23.28 26.18 27.82 28.49
PSNR ↑ Room Counter Kitchen Bonsai
Res x1 x2 x4 x8 x1 x2 x4 x8 x1 x2 x4 x8 x1 x2 x4 x8
PyNeRF 31.73 32.75 32.71 31.43 26.89 27.47 27.72 27.64 30.68 31.45 31.89 30.98 32.74 33.52 33.63 33.08
Zip-NeRF 31.65 32.52 32.78 32.74 28.19 29.15 29.45 29.49 31.72 32.21 32.51 32.53 34.07 34.20 34.03 33.94
Our Model 31.74 32.69 32.86 32.64 28.63 29.51 30.12 30.33 31.17 32.27 33.15 33.07 33.87 34.29 33.97 33.55
SSIM ↑ Bicycle Flowers Garden Stump Treehill
Res x1 x2 x4 x8 x1 x2 x4 x8 x1 x2 x4 x8 x1 x2 x4 x8 x1 x2 x4 x8
PyNeRF 0.710 0.840 0.891 0.896 0.587 0.764 0.836 0.842 0.827 0.887 0.921 0.910 0.745 0.815 0.853 0.843 0.622 0.769 0.845 0.882
Zip-NeRF 0.703 0.834 0.887 0.896 0.565 0.752 0.835 0.866 0.820 0.897 0.923 0.926 0.725 0.807 0.847 0.845 0.591 0.758 0.831 0.874
Our Model 0.716 0.840 0.881 0.890 0.575 0.753 0.829 0.861 0.822 0.890 0.916 0.927 0.726 0.791 0.823 0.827 0.594 0.762 0.845 0.883
SSIM ↑ Room Counter Kitchen Bonsai
Res x1 x2 x4 x8 x1 x2 x4 x8 x1 x2 x4 x8 x1 x2 x4 x8
PyNeRF 0.899 0.917 0.919 0.923 0.845 0.884 0.866 0.875 0.891 0.902 0.917 0.920 0.909 0.935 0.928 0.940
Zip-NeRF 0.910 0.932 0.937 0.936 0.857 0.869 0.875 0.881 0.909 0.924 0.932 0.941 0.946 0.945 0.938 0.939
Our Model 0.907 0.932 0.937 0.937 0.858 0.880 0.897 0.910 0.900 0.912 0.922 0.936 0.939 0.944 0.938 0.941
LPIPS ↓ Bicycle Flowers Garden Stump Treehill
Res x1 x2 x4 x8 x1 x2 x4 x8 x1 x2 x4 x8 x1 x2 x4 x8 x1 x2 x4 x8
PyNeRF 0.211 0.090 0.055 0.059 0.283 0.097 0.060 0.057 0.109 0.058 0.040 0.042 0.191 0.114 0.102 0.094 0.365 0.183 0.106 0.072
Zip-NeRF 0.180 0.106 0.062 0.043 0.306 0.113 0.064 0.047 0.135 0.064 0.044 0.036 0.203 0.121 0.096 0.072 0.390 0.191 0.119 0.071
Our Model 0.217 0.103 0.067 0.052 0.286 0.106 0.071 0.055 0.125 0.059 0.040 0.034 0.228 0.128 0.110 0.096 0.363 0.181 0.098 0.069
@@
LPIPS ↓ Room Counter Kitchen Bonsai
Res x1 x2 x4 x8 x1 x2 x4 x8 x1 x2 x4 x8 x1 x2 x4 x8
Nerfacto 0.320 0.163 0.071 0.093 0.438 0.288 0.166 0.176 0.258 0.113 0.092 0.176 0.233 0.112 0.085 0.160
PyNeRF 0.170 0.066 0.041 0.037 0.206 0.102 0.075 0.066 0.118 0.066 0.043 0.035 0.090 0.050 0.043 0.033
Zip-NeRF 0.141 0.064 0.036 0.024 0.188 0.118 0.064 0.061 0.098 0.064 0.043 0.029 0.079 0.050 0.045 0.037
Our Model 0.151 0.069 0.037 0.025 0.187 0.097 0.054 0.036 0.103 0.060 0.041 0.023 0.086 0.048 0.044 0.036